OpenStudy (imnotgoodatmath):

Evaluate 27 x a^-2 for a =3.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

evaluating a variable expression means what to you @ImNotGoodAtMath ?

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ 27 \times a^{-2} }\). When they say that a=3, plug in 3 instead of a. \(\normalsize\color{black}{ 27 \times 3^{-2} }\) Now solve it, \(\normalsize\color{black}{ 27 \times 3^{-2} =3^3 \times 3^{-2}=3^{3+-2}=3^? }\)

OpenStudy (imnotgoodatmath):

7^2?

OpenStudy (solomonzelman):

Where do you get the ` 7 ` from ?!